OpenStudy (anonymous):

integral of (x+1)/(3x^2)

OpenStudy (anonymous):

Integral ( (1 + 3x)/(1 + x^2) dx ) To solve this, split into two fractions and two integrals. Integral ( 1/(1 + x^2) dx) + Integral ( (3x)/(1 + x^2) dx ) The first integral is easy, because it's a known derivative. It's the derivative of arctan(x). arctan(x) + Integral ( (3x)/(1 + x^2) dx ) The second integral can be solved using substitution. arctan(x) + 3 * Integral ( x/(1 + x^2) dx) arctan(x) + 3 * Integral ( 1/(1 + x^2) x dx) Let u = 1 + x^2. Then du = 2x dx, so (1/2) du = x dx arctan(x) + 3 * Integral ( 1/u (1/2) du ) arctan(x) + 3(1/2) Integral ( 1/u du ) arctan(x) + (3/2) Integral ( 1/u du ) arctan(x) + (3/2) ln|u| + C But u = 1 + x^2, so back-substituting gives us arctan(x) + (3/2) ln|1 + x^2| + C and 1 + x^2 is always positive anyway, so there's no need for the absolute value sign. arctan(x) + (3/2) ln(1 + x^2) + C

OpenStudy (anonymous):

Thankyou Very much. The response was very quick and very helpful.

OpenStudy (anonymous):

Your welcome! Glad I could help!